Fourier cosine integral of f(x) is


A)


cos Ax ff(t)sina tdt da 0


d cos 2x ff(t) cos at dt da


B) cosax ff(t)sintdtdx


D) 25 fcos Ax f f(t)cos At dt da


1
Expert's answer
2021-05-07T14:05:02-0400

"\begin{aligned}
\mathcal{F}(f(x)) &= \sqrt{\frac{2}{\pi}}\int_0^a \cos(at)\cos(\omega t) \,\,\mathrm{d}t\\
&= \frac{1}{\sqrt{2\pi}}\int_0^a \cos((a - \omega)t) + \cos((a + \omega)t) \mathrm{d}t
\\&= \frac{1}{\sqrt{2\pi}}\int_0^a \frac{\sin((a - \omega)t)}{a - \omega} + \frac{\sin((a + \omega)t)}{a + \omega}\mathrm{d}t
\\&= \frac{1}{\sqrt{2\pi}}\frac{\sin((a - \omega)a)}{a - \omega} + \frac{\sin((a + \omega)a)}{a + \omega}\mathrm{d}t
\end{aligned}"